OpenStudy (anonymous):

find the area of the parallelogram A: 38cm2 B: 60cm2 C: 70cm2 D: 168cm2

OpenStudy (anonymous):

Second question find the perimeter of the parallelogram 15.9 ft, 24.84 ft, 31.8 ft, 11.73 ft

OpenStudy (anonymous):

here is the picture for question 1

OpenStudy (anonymous):

here is the picture for question 2

OpenStudy (anonymous):

HELP

Directrix (directrix):

>>the first one i got 70 @l_o_w_key How did you get that?

OpenStudy (anonymous):

2. p=2(a+b) p=2(10.8 + 5.1) P=31.8 does that help?

OpenStudy (anonymous):

@directrix

Directrix (directrix):

Yes, it does. I think the answer to the first question is incorrect. The area of a parallelogram is not equal to the product of its sides.

OpenStudy (anonymous):

i don't understand how i got the first one wrong if you and i both got the same answer..

Directrix (directrix):

We did not get the same answer. I was quoting you. A = b*h A = 12 *5 A = 60 square cm

OpenStudy (anonymous):

oh yeah okay i plugged in the wrong number sorry haha obvi i'm not smart. XD